I just need help setting this up. Ive tried two different ways and neither of them is getting me the right answer.

For the function: f(x) 2x2 + 1, find f(x) f(a) a. 2x+2a revieW T-a f(z +h) -f(z) 蠶 Preview ea syntax error Post this question to forum
Answer #1

b) \frac{f(x + h) - f(x)}{h}

= \frac{f(2x^{2} +1+ h) - f(2x^{2} + 1)}{h}

= (2(2x2 + 1 + h)2 + 1 - 2(2x2 + 1)2 + 1)/h


= (2(4x4 + 1 + h2 + 4x2 + 2h + 4x2h) + 1 - 2(4x4 + 4x2 + 1) + 1)/h

= (8x4+ 2 + 2h2 + 8x2 + 4h + 8x2h + 1 - 8x4 - 8x2 - 2 + 1)/h

= (8x2h + 2h2 + 4h + 2)/h